OpenStudy (anonymous):

Phillip deposited $7,775 into a savings account 14 years ago. The account has an interest rate of 4.5% and the balance is currently $14,398.87. How often does the interest compound?

OpenStudy (ranga):

\[\Large A = P(1 + \frac{ r }{ n })^{nt}\]A = Maturity Amount = $14,398.87 P = Principal = $7,775 r = rate of interest in decimal = 4.5/100 = 0.045 t = time duration of deposit = 14 years n = compounding period Plug in all the values and solve for n.

OpenStudy (anonymous):

okay so A=14,398.87(1+0.045/n)^n14

OpenStudy (ranga):

A=14,398.87 P = $7,775

OpenStudy (anonymous):

OHH crap 14,398.87=7,775(1+0.045/n)^n14

OpenStudy (ranga):

Divide both sides by 7775

OpenStudy (anonymous):

16.85=(1+0.045/n)^n14

OpenStudy (ranga):

Do you have a financial calculator or a TVM table in the book that you can look up?

OpenStudy (anonymous):

no i do not /:

OpenStudy (ranga):

I am trying to search for ways to solve it without those two.

OpenStudy (ranga):

You can take log on both sides and then divide both sides by 14.

OpenStudy (anonymous):

1.2=(1+0.045/n)^n ?

OpenStudy (ranga):

When you take log on both sides it becomes: 1.23 = 14n * log(1+0.045/n) Divide both sides by 14 0.0876 = n * log(1 + 0.045/n) This is a difficult equation to solve without a table to lookup or the proper calculator. But trial and error might work. Try n = 1, 2, 3, 4, etc. and see which one satisfies the equation.

OpenStudy (anonymous):

Okay well thank you for your help!

OpenStudy (ranga):

Wait, how did you get 16.85 = (1+0.045/n)^n14 ?

OpenStudy (ranga):

14,398.87=7,775(1+0.045/n)^(14n) Divide both sides by 7775 1.8519 = (1+0.045/n)^(14n) Take log on both sides log(1.8519) = log{ (1+0.045/n)^(14n) } = 14n * log(1+0.045/n) 0.2676 = 14n * log(1+0.045/n) Divide both sides by 14 0.0191 = n * log(1+0.045/n) Trial and error method with n = 1, 2, 3, etc. Try n = 1 and evaluate the right hand side: 1 * log(1 + 0.045) = log(1.045) = 0.0191 This agrees with our left hand side and we got the answer in the very first attempt! So n = 1 which means the interest is compounded once every year.

OpenStudy (anonymous):

ohhh okay! thank you! :D

OpenStudy (ranga):

you are welcome.

OpenStudy (rockinhood):

Annually!